OpenStudy (anonymous):

Let V be the set of all positive real numbers. Determine whether V isa vector space with the operations below: x+y=xy (additon), cx=x^c (multiplication)

OpenStudy (anonymous):

I think it is, I checked all 10 requirements in the definition of vec space, all satisfied

OpenStudy (anonymous):

i dont think the distribution of scalars over "addition" holds. If c is a scalar in a field, and x and y are vectors, then c(x+y) should equal cx+cy. But the above definition of "addition" doesnt satisfy that. c(x+y)= cxy cx+cy=cxy c^2xy=cxy (false)

OpenStudy (anonymous):

I fond the questions online actually and it does c(x+y)=c(xy)=(xy)^c=x^c y^c = x^c+y^c = cx +cy

OpenStudy (anonymous):

Yes I just wanna point this out